OpenStudy (anonymous):

Shouldn't the derivative of e^(-x) be -xe^(-x-1)? why is it -e^(-x)

OpenStudy (anonymous):

Because you have to mulpiply by the derivative of -x which is -1 You can't use the power rul e like that unless it's a constant number :P .

OpenStudy (anonymous):

You can only use the power rule if the exponent is a constant number NOT a function.

OpenStudy (anonymous):

so e is the constant here right?

OpenStudy (anonymous):

yes. e is 2.7 something something something ... i think

OpenStudy (anonymous):

so you can never have x as an exponent to use the chain rule?

OpenStudy (anonymous):

even say (x+1) as an exponent?

OpenStudy (anonymous):

to use the power rule. Chain rule is not the power rule.

OpenStudy (anonymous):

sorry sorry. i meant power rule

OpenStudy (anonymous):

right. just like using the power rule for \[y=10^x\] would be inappropriate

OpenStudy (perl):

e^u = e^u * u'

OpenStudy (perl):

where u is more complicated than just x

OpenStudy (anonymous):

@perl so the derivative of e^(x+1) would be e^(x+1) * (1)?

OpenStudy (callisto):

Recall:\[\frac{d}{dx}e^x=e^x\] First example: \[\frac{d}{dx}e^{-x}=e^{-x}\frac{d}{dx}(-x) = -e^{-x}\] Second example: \[\frac{d}{dx}e^{x+1}=e^{x+1}\frac{d}{dx}(x+1) = e^{x+1}\] You can use substitution, if that makes it easier to understand. let u = x+1 \[\frac{d}{dx}e^{x+1}=\frac{d}{du}e^u \times \frac{d}{dx}(x+1)= e^{u} \times 1 = e^{x+1}\]

OpenStudy (anonymous):

wait. let me process it

OpenStudy (anonymous):

ok. i think i get it now. it's just a rule that would in my opinion be better to learn than understand anyway

OpenStudy (anonymous):

and if it was e^(-x+1) would it be -e^(-x+1)?

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

Yes :) How do you get that?

OpenStudy (anonymous):

i just learnt it! if the power is negative then the constants sign will also change, but the exponent will stay the same

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

and i wont be able to use power rule even if it is e^(2x)

OpenStudy (anonymous):

?

OpenStudy (callisto):

You can actually use chain rule to get that. \[y=e^{-x+1}\] Let u = -x+1 \[\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx} = \frac{d}{du}(e^u) \times \frac{d}{dx}(-x+1)=-e^u = -e^{-x+1}\]

OpenStudy (anonymous):

The negative sign comes from the fact that \[\frac{d}{dx} -x = -1\] if for instance the question was \[ \frac{d}{dx} e^{-2x+1}\] the answer would then be \[-2e^{-2x+1}\] because -2 is the derivative of -2x+1 if for instance the question was \[ \frac{d}{dx} e^{-5x+1}\] the answer would then be \[-5e^{-2x+1}\] because -5 is the derivative of -5x+1

OpenStudy (callisto):

You're actually using chain rule instead of power rule in doing these examples. For e^(2x) \[y = e^{2x}\] Let u = 2x \[\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx} = \frac{d}{du}(e^u) \times \frac{d}{dx}(2x)=...?\]

OpenStudy (anonymous):

@freewilly922 for e^(-5x+1) shouldn't it be -5e^(-5x+1)? you have -5e^(-2x+1)

OpenStudy (anonymous):

yes it should I was copying and pasting and messed up. Good catch. And sorry.

OpenStudy (anonymous):

@Callisto the derivative of 2x is 2. so should it be 2e^(2x)?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

good work @freewilly922 very good explanation.

OpenStudy (anonymous):

YIPPEEE!!

OpenStudy (anonymous):

i wish i could give 2 best responses

OpenStudy (anonymous):

freewilly needs it more i guess

OpenStudy (anonymous):

@Callisto was a better response.

OpenStudy (anonymous):

u both did great. thank you very much @Callisto and @freewilly922